OpenStudy (anonymous):

if f(x)= x^4 - 4x evaluate lim x->-1 (f(x)-f(-1))/(x+1) a. -8 b. 0 c. 1 d. 2 e. 4 PLEASE HELP!

OpenStudy (anonymous):

it is -8. By first principle,f'(-1)=lim h->0 (f[-1+h]-f[-1])/h

OpenStudy (anonymous):

this is \[\lim_{x\rightarrow -1}\frac{f(x)-f(-1)}{x+1}\]?

OpenStudy (anonymous):

\[f(-1)=5\] so you get \[\frac{x^4-4x-5}{x+1}=\frac{(x+1) (x^3-x^2+x-5)}{x+1}\]

OpenStudy (anonymous):

cancel and replace x b y -1 to get your answer

OpenStudy (anonymous):

steps clear?

OpenStudy (anonymous):

are you sure thats factored right??